OpenStudy (anonymous):

Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1). y = -4x + 1 y = -2x + 9 y = 4x + 1 y = -1/2 x + 9

OpenStudy (anonymous):

First find the slope between the two points

OpenStudy (anonymous):

\[\Large slope=m=\frac{ y_2-y_1 }{ x_2-x_1}\]

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

for the slope @doulikepiecauseidont

OpenStudy (anonymous):

OK, now we have an equation called point-slope formula. All you need is a point on the graph (we have two so we can pick one) and we also need the slope of the linewhich we found to be -2 \[\Large y-y_1=m(x-x_1)\] Where y & x don't change (x_1, y_1) are one of the points m=slope

OpenStudy (anonymous):

so y = -2x + 9

OpenStudy (anonymous):

Correct, good job!